OpenStudy (anonymous):

during the first part of a trip a canoeist travels 56 miles at a certain speed. the canoeist travels 22 miles on the second trip at a speed 5 mph slower. the total time for the trip is 3 hrs. what was the speed on each part of the trip?

OpenStudy (radar):

d=st where distance equal speed times time. The time for the complete trip is 3 hrs. Trip was in two parts where the time for each part would be: 56/speed for first leg 22/speed for second leg Let s = first leg speed s-5 would be speed of second leg (given) then we have this equation 56/s + 22/(s-5) = 3 hrs. The rest is algebra

OpenStudy (radar):

First using the LCD of s(s-5) add the left hand terms (78s-280)/(s^2-5s)=3 Now cross multiply getting: \[3s ^{2}-15s=78s-280\]Consolidate and put in standard form:\[3s ^{2}-93s+280=0\]

OpenStudy (radar):

Solve using the quadratic formula for speed.

OpenStudy (radar):

Using calculator I get speed of 27.62 for first leg and 22.62 for second leg. Warning, something doesn't seem right, that is awful fast for paddling a canoe lol. Check everything.

OpenStudy (radar):

@dumbcow, did I make an error?

OpenStudy (radar):

Sometimes the data furnished is not logical for the problem, this may be the case, plus they must of paddled fast!

OpenStudy (radar):

@Mertsj could you, at your leisure (of course) verify this solution?